OpenStudy (anonymous):

use f(x)= x/ absolute value of x, to find lim x> infinity, lim x> - infinity, and all horizontal asymptotes??

OpenStudy (anonymous):

@CarlosGP, would lim f(x) x>- infinity = -1 ? and lim f(x) x->infinity = 1 ??

OpenStudy (anonymous):

@Hero ?

hero (hero):

Actually, for this one, it is tricky.

hero (hero):

Simply because.. \[\frac{x}{x} = 1\] and \[-\frac{x}{x} = -1\]

hero (hero):

So basically, when that happens, it means there's no horizontal asymptote.

hero (hero):

If yo graph it, you'll see two lines where f(x) = 1 and f(x) = -1 graphically

hero (hero):

But there's no horizontal asymptote

OpenStudy (anonymous):

hmmm, yeah I guess I can see why there wouldn't be a horizontal asymptote !

OpenStudy (anonymous):

(:

OpenStudy (anonymous):

think you could help me out with 2 more? D: (:

hero (hero):

That's the reason why I try to simplify these first before attempting to find asymptotes.

OpenStudy (anonymous):

i wasn't taught how to simplify them to find the asymptotes, we just graph :o

OpenStudy (anonymous):

find the verticle asymptote of f(x)= 1/(x^2 -4)

OpenStudy (anonymous):

and describe the behavior of f(x) to the left and right of each asymptote